OpenStudy (anonymous):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can make as many corrections as she wants. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can receive. Identify this as converging or diverging.

OpenStudy (anonymous):

She lost 35 points She can get 3/10 back 35*3/10=105/10=10.5 points maximum for new grade of 75.5 maximum. converging on 75.5 or 10.5 depending on your viewpoint 65+.3*n=s

OpenStudy (anonymous):

I figured out that she can get a maximum score of 75.5 but what do i do next? It asks me to create an equation for the sum of the geometric series. I'm confused about which values to plug into the equation. @jaida.115

OpenStudy (bradely):

a = 65 r= 3/10 sn= a(1-r^n)/(1-r) =65(1-(3/10)^n)/(1-(3/10)) u can simplify u get source: http://www.mathskey.com/question2answer/ask?cat=3

OpenStudy (anonymous):

thanks @bradely

OpenStudy (anonymous):

What do I have to substitute for n? @bradely